OpenStudy (anonymous):

convert to mixed number 12/5

OpenStudy (anonymous):

12/5=(10+2)/5=2(2/5)

OpenStudy (anonymous):

Divide 12 by 5. You get 2 and have 2 left over. Just put the remainder as the numerator. So it's 2 2/5.

OpenStudy (lgbasallote):